6. A individual is borrowing 145,000 for a 15 year loan at 3.57% per year compounded monthly. Calculate her monthly payment. 7. For the loan described under Q6, what would be the amount owed to the bank if the individual decides to sell the house immediately after the 110h mortgage payment?
Answer #1

6 . Principal =145,000

Rate=3.57%

Time=1 month

=1 / 12

Interest =P*R*T/ 100

=145000*3.57*1/12 whole divided by 100

=143.125

Monthly payment=143.125

7. Question is incomplete.